# DrivingGen: A Comprehensive Benchmark for Generative Video World Models in Autonomous Driving

### v1

@@ A.2 Benchmarks for evaluating generative world models
The rapid progress of open- and closed-source video generation has driven the creation of many benchmarks (huang2023vbenchcomprehensivebenchmarksuite; huang2024vbenchcomprehensiveversatilebenchmark; bansal2024videophyevaluatingphysicalcommonsense; ning2023videobenchcomprehensivebenchmarktoolkit; liao2024evaluationtexttovideogenerationmodels; fan2024aigcbenchcomprehensiveevaluationimagetovideo; wang2024mattersnovelvisualphysicsbased), such as VBench, which evaluates models with multifaceted metrics based on human-collected prompts. Recently, evaluations have expanded to open, dynamic, and complex world-simulation scenarios (yue2025ewmbenchevaluatingscenemotion; duan2025worldscoreunifiedevaluationbenchmark; li2025driversenavigationworldmodel; qin2024worldsimbenchvideogenerationmodels; kwon2025stableworldmodelsmeasuring). WorldScore (duan2025worldscoreunifiedevaluationbenchmark) measures generated videos using explicit camera trajectory layouts. However, a comprehensive driving-world benchmark is still lacking due to limited test sample diversity, heterogeneous input modalities, and the absence of driving-specific metrics. Recent works (gao2024vista; hassan2024gemgeneralizableegovisionmultimodal) mainly adopt Frechet Video Distance (FVD) and Average Displacement Error (ADE) for trajectory alignment, while GEM (hassan2024gemgeneralizableegovisionmultimodal) adds human video evaluations that are subjective and hard to scale. The closest effort, ACT-Bench (arai2024actbenchactioncontrollableworld), focuses solely on trajectory alignment and overlooks key aspects such as video and trajectory distribution, quality, and temporal consistency.
@@ B.2 Details of Our SLAM Pipeline and compariision with Others
Comparison with Other SLAM Pipelines. We evaluated our reconstruction pipeline against those used in recent driving world-model systems. Concretely, we compare the successful reconstruction rate and trajectory accuracy (ADE) on 20 nuPlan videos generated with Vista from our early experiments. A run is counted as successful if the SLAM system returns a valid camera trajectory without numerical failure. The results are summarized in Table 4. Compared to the GEM pipeline (DROID-SLAM (teed2022droidslamdeepvisualslam) + Depth-Anything v2 (yang2024depthv2)) and the DrivinDojo pipeline (COLMAP (schoenberger2016mvs; schoenberger2016sfm) with scale aligned to ground truth), our basic version (Ours w/o failure handling) achieves a similar successful reconstruction rate (17/20 vs. 17/20 and 16/20) and a comparable ADE (15.18 vs. 14.61 and 14.99). When we enable our failure-handling strategy (Ours w/ failure handling), the successful rate increases to 20/20, while the ADE remains in the same ballpark (16.84). This trade-off is important for DrivingGen: the benchmark needs robust reconstruction on all videos rather than dropping harder cases and evaluating on a subset of “easy” videos. Overall, our SLAM pipeline is more robust than existing pipelines by handling reconstruction failure explicitly.
@@ B.3 Fréchet Trajectory Distance (FTD)
Representation model and input. We use MTR’s agent_polyline_encoder $\phi(\cdot)$ (shi2023motiontransformerglobalintention). Crucially, MTR consumes a fixed temporal horizon $H$.
@@ B.4 Objective Image Quality
Motivation and background. Pulse–width modulation (PWM) in vehicle lighting and roadside luminaires induces temporal luminance modulation that, when sampled by rolling-shutter cameras, can alias into low-frequency flicker and degrade detection and tracking. The IEEE Automotive P2020 standard formalizes Modulation Mitigation Probability (MMP) to quantify whether such modulation is sufficiently suppressed during operation (ieee2018ieee; 9969530). We implement MMP on the frame-mean luminance to provide a robust and efficient evaluation signal.

### v2

@@ A.2 Benchmarks for evaluating generative world models
The rapid progress of open- and closed-source video generation has driven the creation of many benchmarks (Huang et al., 2023; 2024; Bansal et al., 2024; Ning et al., 2023; Liao et al., 2024; Fan et al., 2024; Wang et al., 2024g), such as VBench, which evaluates models with multifaceted metrics based on human-collected prompts. Recently, evaluations have expanded to open, dynamic, and complex world-simulation scenarios (Yue et al., 2025; Duan et al., 2025; Li et al., 2025d; Qin et al., 2024; Kwon et al., 2025). WorldScore (Duan et al., 2025) measures generated videos using explicit camera trajectory layouts. However, a comprehensive driving-world benchmark is still lacking due to limited test sample diversity, heterogeneous input modalities, and the absence of driving-specific metrics. Recent works (Gao et al., 2024; Hassan et al., 2024) mainly adopt Frechet Video Distance (FVD) and Average Displacement Error (ADE) for trajectory alignment, while GEM (Hassan et al., 2024) adds human video evaluations that are subjective and hard to scale. The closest effort, ACT-Bench (Arai et al., 2024), focuses solely on trajectory alignment and overlooks key aspects such as video and trajectory distribution, quality, and temporal consistency.
@@ B.2 Details of Our SLAM Pipeline and compariision with Others
Comparison with Other SLAM Pipelines. We evaluated our reconstruction pipeline against those used in recent driving world-model systems. Concretely, we compare the successful reconstruction rate and trajectory accuracy (ADE) on 20 nuPlan videos generated with Vista from our early experiments. A run is counted as successful if the SLAM system returns a valid camera trajectory without numerical failure. The results are summarized in Table 4. Compared to the GEM pipeline (DROID-SLAM (Teed & Deng, 2022) + Depth-Anything v2 (Yang et al., 2024c)) and the DrivinDojo pipeline (COLMAP (Schönberger et al., 2016; Schönberger & Frahm, 2016) with scale aligned to ground truth), our basic version (Ours w/o failure handling) achieves a similar successful reconstruction rate (17/20 vs. 17/20 and 16/20) and a comparable ADE (15.18 vs. 14.61 and 14.99). When we enable our failure-handling strategy (Ours w/ failure handling), the successful rate increases to 20/20, while the ADE remains in the same ballpark (16.84). This trade-off is important for DrivingGen: the benchmark needs robust reconstruction on all videos rather than dropping harder cases and evaluating on a subset of “easy” videos. Overall, our SLAM pipeline is more robust than existing pipelines by handling reconstruction failure explicitly.
@@ B.3 Fréchet Trajectory Distance (FTD)
Representation model and input. We use MTR’s agent_polyline_encoder $\phi(\cdot)$ (Shi et al., 2023). Crucially, MTR consumes a fixed temporal horizon $H$.
@@ B.4 Objective Image Quality
Motivation and background. Pulse–width modulation (PWM) in vehicle lighting and roadside luminaires induces temporal luminance modulation that, when sampled by rolling-shutter cameras, can alias into low-frequency flicker and degrade detection and tracking. The IEEE Automotive P2020 standard formalizes Modulation Mitigation Probability (MMP) to quantify whether such modulation is sufficiently suppressed during operation (IEEE P2020 Working Group, 2018; IEEE, 2022). We implement MMP on the frame-mean luminance to provide a robust and efficient evaluation signal.